OpenStudy (anonymous):

Can the set of lengths be the side lengths of a right triangle? 8 ft, 15 ft, 17 ft

OpenStudy (anonymous):

Yup! use the pythagorean theorem.. a^2+b^2=c^2 64+225=289

OpenStudy (anonymous):

8² + 15² = 17² easier to see/visualize like this :P the first two numbers when squared have to equal the last number squared